OpenStudy (mony01):

find 10 partial sums of the series, graph sequence of partial sum and sequence of terms, is it convergent or divergent?

OpenStudy (mony01):

\[\sum_{n=2}^{\infty}\frac{ 1 }{ n(n+2) }\]

OpenStudy (usukidoll):

dang if only I still remember... I do know that there are a lot of tests and each of them have their own unique rules to determine convergent or divergent...

OpenStudy (mony01):

you dont know what test i can use to solve this?

OpenStudy (anonymous):

comparison and P-series

OpenStudy (anonymous):

1/(n(n+2)) = 1/(n^2 + 2n) < 1/(n^2) 1/n^2 is convergent by P-series, hence 1/(n(n+2)) is convergent by comparison

OpenStudy (anonymous):

as far as partial sum goes, just grab a calculator

OpenStudy (anonymous):

or use wolfram alpha

OpenStudy (mony01):

i know the answer but since its convergent how would i figure out the answer

OpenStudy (anonymous):

well it ask for the 10th partial sum, meaning the sum of the first 10 term. Since n start at 2, the tenth term is n = 11

OpenStudy (mony01):

so i add all the sums?

OpenStudy (anonymous):

for the value of the 10th partial so, yes. For the sequence of the partial sums, you need find 1st sum, 2nd sum, ... , 10th sum, and plot those values

OpenStudy (anonymous):

10th partial *sum*

OpenStudy (mony01):

i did a table and plot those points, but dont i need another graph?

OpenStudy (anonymous):

remember, partial sum isn't the same as the value of the sequence at particular index

OpenStudy (usukidoll):

geez don't have to be rude... why don't you I don't know just distribute the n on the denominator and use p-series test. any exponent greater than 1 converges . have a nice day *sarcasm*

OpenStudy (mony01):

@sourwing i dont quite understand this?

OpenStudy (usukidoll):

D:

OpenStudy (mony01):

@UsukiDoll who are you referring with that comment?

OpenStudy (usukidoll):

you

OpenStudy (anonymous):

what graph did you make?

OpenStudy (mony01):

what did i do? @UsukiDoll

OpenStudy (usukidoll):

you said "you dont know what test i can use to solve this?" It's been a while for me k... chill and sourwing gave the answer already... use p series it's much easier just distribute the n my gawd!

OpenStudy (mony01):